OpenStudy (anonymous):

find the area of the region described by y≥2|x-3|+4 and x-2y≥-20

OpenStudy (anonymous):

looks like some sort of triangle http://www.wolframalpha.com/input/?i=+y%E2%89%A52 |x-3|%2B4+and+x-2y%E2%89%A5-20

OpenStudy (anonymous):

x-3|%2B4+and+x-2y%E2%89%A5-20 what does this mean?

OpenStudy (freckles):

I think it was all one link

OpenStudy (freckles):

i think you start copying from the h to the last 0 there it is suppose to be a link if i'm not mistaken

OpenStudy (freckles):

Anyways we could try to draw this ourselves.

OpenStudy (freckles):

Let's look at the easiest part first. How do you graph x-2y=-20?

OpenStudy (anonymous):

y≥-x/(-2)-(-20)/(-2)

OpenStudy (anonymous):

do you question marks in some diamond shape? I dont know if you see that, but i dont know how to make it stop

OpenStudy (freckles):

i have to keep refreshing the page for it to unshow

OpenStudy (anonymous):

oh ok is this right? y≥-x/(-2)-(-20)/(-2)

OpenStudy (freckles):

\[x-2y \ge -20 \] So I guess you subtracted x on both sides \[-2y \ge -x-20\] you should divide by the -2 on both sides (when you divide both sides by a negative number remember to flip the inequality) \[y \le \frac{x}{2}+10\]

OpenStudy (freckles):

well the direction of the inequality you have is not eating the right part and you have to many negative over that on that 20/2 part

OpenStudy (freckles):

\[x-2y \ge -20 \\ -2y \ge -20-x \\ -2y \ge -x-20 \\ y \le \frac{-x}{-2}-\frac{20}{-2}\]

OpenStudy (freckles):

I think this is what you meant to write

OpenStudy (anonymous):

oh ok i see my mistake

OpenStudy (freckles):

and you can cancel those negative's out in the x/2 term and the same for that 20/2 part \[y \le \frac{x}{2}+10 \]

OpenStudy (anonymous):

ok

OpenStudy (freckles):

ok we have that where we can identify the slope and y-intercept of y=x/2+10 now

OpenStudy (freckles):

we could go ahead and look at the absolute value part so we have y=2|x-3|+4 if x-3>0 then |x-3|=x-3 if x-3<0 then |x-3|=-(x-3) we are going to have to write y=2|x-3|+4 as a piecewise function

OpenStudy (anonymous):

ok

OpenStudy (freckles):

or better yet if you know about translating graphs then we can skip that

OpenStudy (freckles):

do you know the vertex of y=2|x-3|+4?

OpenStudy (anonymous):

y≥2|x|7 ??

OpenStudy (freckles):

? the vertex of say y=a|x-h|+k is (h,k) if a>0 then the absolute value opens up if a<0 then the absolute value opens down

OpenStudy (freckles):

so what is the vertex of y=2|x-3|+4 try to use what I just said to answer

OpenStudy (anonymous):

...i dont know...

OpenStudy (anonymous):

sorry

OpenStudy (freckles):

ok so do you know that |x-3|=x-3 when x-3>0 and |x-3|=-(x-3) when x-3<0 because you know this then I guess we can go back to the piecewise thing

OpenStudy (anonymous):

ok

OpenStudy (freckles):

so can you write y=2|x-3|+4 as a piecewise function?

OpenStudy (anonymous):

y=2x-3+4 ?

OpenStudy (freckles):

remember |x-3|=x-3 if x-3>0 |x-3|=-(x-3) if x-3<0

OpenStudy (freckles):

do you know calculus? We can use that to find the area. I was really hoping to draw the graph though but I don't think you can do that without knowing how to write absolute value function as a piecewise function or using the fact that the vertex is (3,4) and the absolute value function is open up because 2>0 I guess you can plug in values for x to see what y is (I hate this approach though because you have to plot enough points to get an almost accurate graph)

OpenStudy (anonymous):

its ok i will just try to figure it our myself...thanks though :)